What is radius of curvature in projectile motion explain
 

Grade:11

1 Answers

Vikas TU
7 years ago
the radius of curvature is defined at each point, given by the formula faust9 gave. The radius of curvature is given at the nozzle and the question asks for the radius of curvature at the vertex of the parabola. JohnDubYa's point that this is a parabola is important: set up a coordinate system so that the nozzle is at (0,0) and and the y axis is parallel to the axis of the parabola. You can write the parabola as y= ax2+ bx. y'= 2ax+ b which, at x= 0, is y'= b. You can find b from the "3- 4- 5" information. y"= 2a and you find a from the curvature at x= 0. Once you know the equation of the parabola you can use the fact y= -(g/2)t2+ vyt and x= vxt to find vx and vy[\sub] (the components of velocity at t= 0) to find the initial velocity